4. Hai Ying is a junior at CSI who has applied for two summer internship programs. Since she has excellent grades in statistics, Hai Ying feels that she has a 70% chance of being offered an internship at the Prestigious Financial Firm and an 80% chance of being offered an internship at the Widely Respected Non-profit. If she receives an offer from the Non-profit, she believes that she has an 85% chance of receiving an offer from the Financial Firm. a. Calculate the probability that both the Financial Firm and the Non-profit will make her an offer. b. Calculate the probability that she will be offered at least one of the internships. c. If she does not receive an offer from the Financial Firm, calculate the probability that she will not receive an offer from the Non-profit either.
Answer #1

a)P(both make an offer) =P(receive offere from Non-profit )*P(receive from Financial Firm |non profit)

=0.80*0.85 =0.68

b)

she will be offered at least one of the internships =P(financial)+P(non profit)-P(both)=0.7+0.8-0.68=0.82

c)

P(not from Non profit |not from Financial) =(1-0.82)/(1-0.70)=0.60